# The FineWeb Datasets: Decanting the Web for the Finest Text Data at Scale

### v1

@@ 2 Background
Given that the focus of our work is to carefully design an effective Common Crawl-based pretraining dataset, we now briefly discuss the filtering and deduplication used in past public datasets. OSCAR [25] processes Common Crawl using a pipeline inspired by that of Touvron et al. [2], which uses a fastText-based language classifier [26] to filter pages based on their language and then performs deduplication at the line level using a non-cryptographic hash algorithm. C4 [15] uses langdetect [27] to filter out non-English pages, then applies a series of heuristic filters (retaining only those lines that end in a terminal punctuation mark, removing short lines, discarding any page that contains a word from a “bad words” list, etc.), and finally deduplicates over three-line windows. CC100 [28] uses the cc_net pipeline, which includes fastText for language identification, performs paragraph-level hash-based deduplication, and retains only the text that is assigned a low perplexity by a n-gram language model trained on Wikipedia. The Pile [29] is a composite dataset that includes “Pile-CC”, a collection of text from one Common Crawl snapshot that uses pycld2 [30] for language detection, jusText [31] for boilerplate removal, a classifier to filter out pages that are dissimilar form WebText (described below), and fuzzy deduplication using MinHash [32]. ROOTS [33] includes text from the pre-processed web crawl OSCAR with additional heuristic-based filtering and SimHash-based deduplication. RedPajama [34] is a composite dataset that includes Common Crawl-sourced text processed using the cc_net pipeline as well as quality filtering using a classifier trained to distinguish Wikipedia level content from random Common Crawl samples. SlimPajama [35] further processed RedPajama by removing short documents and performing additional fuzzy MinHash-based deduplication. RefinedWeb [36] uses trafilatura [37] for text extraction, fastText for language identification, heuristic rules inspired by MassiveText (discussed below) to filter data, and both MinHash (fuzzy) and ExactSubstr (exact) deduplication. RedPajama v2 [34] has 84 Common Crawl snapshots released unfiltered and non-deduplicated but with labels from filtering techniques from cc_net, C4, MassiveText, RefinedWeb and others, as well as deduplication labels for exact (Bloom filter) and fuzzy (MinHash) deduplication. Finally, Dolma [38] has a Common Crawl-based subset that uses fastText for language classification, heuristic rules from MassiveText and C4 for quality filtering, rules- and classifier-based toxicity filtering, and URL, document and paragraph-level deduplication using a Bloom filter.
@@ 3 Building FineWeb
Our design of FineWeb is primarily empirical; i.e., we performed a series of “data ablation” experiments to test different methods at each stage of the pipeline. In this section, we chronicle our experimental results and design choices. Our results are fully reproducible with code in our datatrove repository.
@@ 3.1 Experimental setup
All training was performed using the nanotron library. Data ablation models all had 1.82B parameters (including embeddings), used the Llama architecture [2] with a sequence length of 2048, a global batch size of ~2 million tokens, and the GPT2 tokenizer [20]. Within a given experiment, all models were trained on the same amount of data for the same number of steps. Filtering ablations were trained on ~28 billion tokens (roughly the Chinchilla-optimal training size for this model size [1]), while some deduplication ablations and runs to confirm cumulative relative performance improvements after each step of filtering were conducted on 350 billion tokens. We make all models trained for our ablations publicly available on our dataset repository. In total, we trained over 70 models on our internal cluster, for an estimated total of 80,000 H100 GPU hours.
@@ 3.2 Text extraction
Common Crawl data is available in two different formats: WARC and WET. WARC (Web ARChive format) files contain the raw data from the crawl, including the full page HTML and request metadata. WET (WARC Encapsulated Text) files provide a text-only version of crawled websites by using htmlparser [48]. While WET files are commonly used as a starting point for dataset creation, similarly to Gao et al. [29], we found that WET files retained too much boilerplate and menu text. We therefore experimented with extracting the text content from the WARC files using the open source trafilatura library [49], which from visual inspection of the results provided good quality extraction when compared to other available libraries (less boilerplate and menu text). Custom text extraction is relatively costly, but its effects are felt on model performance: Fig. 2 shows the performance of ablation models trained on either trafilatura applied to WARC data or WET data, with minimal additional filtering (fastText language identification to filter samples with English as the highest probability label). Using trafilatura-extracted text from WARC files clearly results in a more performant model and we therefore use WARC-based data in all of our following experiments.
Figure 1: Trafilatura-extracted WARC vs WET 28B tokens ablation study. Custom text extraction outperforms the default WET data.
@@ 3.4 Deduplication
Following RefinedWeb [50], we experimented with MinHash, a fuzzy hash-based deduplication technique that scales efficiently to many CPU nodes and allows tuning of similarity thresholds (by controlling the number and the number of hashes per bucket) as well as the length of the subsequences considered (by controlling the n-gram size). We chose to collect each document’s 5-grams, obtained using an English word tokenizer [57], and computed MinHashes using 112 hash functions in total, split into 14 buckets of 8 hashes each — targeting documents that are at least 75% similar. Documents with the same 8 MinHashes in any bucket are considered duplicates of each other. We further discuss deduplication parameters in the Appendix.
@@
To attempt to improve upon individual deduplication of each snapshot, we additionally experimented with “lighter” global deduplication techniques. Ultimately, none of these techniques improved performance over independent per-snapshot deduplication. A full description of these methods and results are in the Appendix.
@@ 3.6 Developing additional heuristic filters
Past work has mainly developed heuristic filters through data inspection [15]. In this work we devised a more systematic process for designing heuristic filters and tuning their thresholds. We started by collecting over 50 high-level statistics ranging from document-level metrics (e.g. number of lines, avg. line/word length, etc) to inter-document repetition metrics (inspired by MassiveText [1]) on both a high- and low-quality web dataset. Specifically, we used the individually and globally deduplicated versions of the 2013-48 snapshot (previously mentioned in Section 3.4) as our “high-quality” and “low-quality” datasets respectively. We then identified metrics for which the distribution of values differed significantly across the two datasets, inspected the histograms of the two distributions and empirically chose thresholds that would target sections of the histogram where the lower quality dataset frequency was higher than on the corresponding higher quality dataset section. As an example, we plot the distribution of the fraction of lines ending with punctuation metric in  Fig. 8. We can see that the higher quality dataset has in general higher document density for larger values of our metric, and, in particular, the lower quality dataset has a much higher density of documents for values < 0.12. We thus conclude that documents with a fraction of lines ending with punctuation < 0.12 are generally lower quality and use this value as a tentative threshold to filter documents. Following this process for all metrics yielded 17 candidate metric-threshold pairs.
@@
We then assessed the effectiveness of these 17 newly created filters by conducting several 28B token ablation runs on the 2019-18 crawl. Out of all those runs, we identified three filters (see their ablations runs in Fig. 7) that demonstrated the most significant improvements on the aggregate benchmark score. Specifically, the chosen filters remove documents where the fraction of lines ending with punctuation is $<=0.12$ (10.14% of tokens removed vs. 30% from the original C4 terminal punctuation filter), where the fraction of characters in duplicated lines is $>=0.1$ (12.47% of tokens removed; the original MassiveText threshold for this ratio is $>=0.2$), and/or where the fraction of lines shorter than 30 characters is $>=0.67$ (3.73% of tokens removed). When applying the three together, ~22% of tokens were removed and the aggregate score increased by about 1% in the 28B token ablations. These filters allowed us to further improve performance and, notably, surpass the C4 dataset performance while filtering out a smaller proportion of data.
@@ 3.7 The final FineWeb dataset
In Fig. 10 we compare FineWeb with the following commonly used openly accessible web-scale datasets: RefinedWeb (500B tokens) [50], C4 (172B tokens) [15]; the Common Crawl-based part of Dolma 1.6 (3T tokens) and 1.7 (1.2T tokens) [58], The Pile (340B tokens) [29], SlimPajama (627B tokens) [35], the deduplicated variant of RedPajama2 (20T tokens) [34], English CommonCrawl section of Matrix (1.3T tokens) [59], English CC-100 (70B tokens) [60], and Colossal-OSCAR (850B tokens) [61]. Notably, FineWeb shows strong performance and FineWeb-Edu (detailed below) outperforms all other open datasets we compared on our aggregate group of tasks, further validating the design choices we made.
@@ 4 FineWeb-Edu
To build the synthetic annotations, we use Llama-3-70B-Instruct to score 460,000 samples from FineWeb for their educational quality on a scale from 0 to 5. We explored several prompt formats to automatically extract an educational score using an LLM and found that the additive scale used in previous work Yuan et al. [63] worked best. It allows the LLM to evaluate each criterion and build the score step-by-step, unlike the single-rating scale [64] which assigns a fixed score based on predefined categories. To avoid having the LLM favor highly technical pages like arXiv abstracts and submissions, we prompted it to focus on grade-school and middle-school level knowledge. The prompt used for synthetic annotations is in the Appendix.
@@
To confirm the effectiveness of education filtering at a larger scale, we conducted a larger ablation training a 1.82B model on 350 billion tokens, similar to the FineWeb filtering ablations mentioned above. As shown in Fig. 10 and Fig. 11, we observed that FineWeb-Edu surpasses FineWeb and all other open web datasets, with quite remarkable improvements on educational benchmarks such as MMLU, ARC and OpenBookQA. Specifically, MMLU score increases from 33% to 37%, a relative improvement of approximately 12%, and ARC score goes from 46% to 57%, an improvement of about 24%.
On MMLU, FineWeb-Edu can match the final performance of Matrix with almost 10x fewer tokens, demonstrating the effectiveness of classifiers trained on LLM annotations for large-scale data filtering. Additional evaluation plots can be found in the Appendix.
@@ 5 Bias analyses
Language models are known to reflect the biases present in their pretraining datasts [67, 68, 69, 70, 71, 72]. To provide a brief picture of dataset bias in FineWeb and FineWeb-Edu, we focus on subgroups recognised as “sensitive” or “protected” in English-speaking countries. These are a subset of subgroups that are historically subject to discrimination and are disproportionately the target of negative societal norms such as stereotyping, which is reflected in text-based language consumed for a dataset. We find that the FineWeb dataset has a relative overrepresentation of words that reflect hegemonic norms, known to be overrepresented in online text [67], such as ‘man’ and ‘christian’. Although biases across the gender, religion, and age subgroups examined are not strong, we see the most skewed association between religion words and intimacy, such as ‘christian dating’ and ‘jewish singles’. Fittingly, the FineWeb-Edu dataset captures associations that are less tied to intimacy compared to FineWeb and more expected from educational content of history and health, such as ‘man’ being associated to ‘king’, and ‘woman’ associated to ‘pregnancy’. Further details are provided in the Appendix.
@@ Appendix C Linked resources
Appendix D Deduplication
D.1 Deduplication parameters
@@
D.2 Measuring the effect of deduplication
@@
D.3 Alternative global deduplication

### v2

@@ 2 Background
Given that the focus of our work is to carefully design an effective Common Crawl-based pretraining dataset, we now briefly discuss the filtering and deduplication used in past public datasets. OSCAR [25] processes Common Crawl using a pipeline inspired by that of Touvron et al. [2], which uses a fastText-based language classifier [26] to filter pages based on their language and then performs deduplication at the line level using a non-cryptographic hash algorithm. C4 [15] uses langdetect [27] to filter out non-English pages, then applies a series of heuristic filters (retaining only those lines that end in a terminal punctuation mark, removing short lines, discarding any page that contains a word from a “bad words” list, etc.), and finally deduplicates over three-line windows. CC-100 [28] uses the cc_net pipeline, which includes fastText for language identification, performs paragraph-level hash-based deduplication, and retains only the text that is assigned a low perplexity by a n-gram language model trained on Wikipedia. The Pile [29] is a composite dataset that includes “Pile-CC”, a collection of text from one Common Crawl snapshot that uses pycld2 [30] for language detection, jusText [31] for boilerplate removal, a classifier to filter out pages that are dissimilar form WebText (described below), and fuzzy deduplication using MinHash [32]. ROOTS [33] includes text from the pre-processed web crawl OSCAR with additional heuristic-based filtering and SimHash-based deduplication. RedPajama [34] is a composite dataset that includes Common Crawl-sourced text processed using the cc_net pipeline as well as quality filtering using a classifier trained to distinguish Wikipedia level content from random Common Crawl samples. SlimPajama [35] further processed RedPajama by removing short documents and performing additional fuzzy MinHash-based deduplication. RefinedWeb [36] uses trafilatura [37] for text extraction, fastText for language identification, heuristic rules inspired by MassiveText (discussed below) to filter data, and both MinHash (fuzzy) and ExactSubstr (exact) deduplication. RedPajama v2 [34] has 84 Common Crawl snapshots released unfiltered and non-deduplicated but with labels from filtering techniques from cc_net, C4, MassiveText, RefinedWeb and others, as well as deduplication labels for exact (Bloom filter) and fuzzy (MinHash) deduplication. Finally, Dolma [38] has a Common Crawl-based subset that uses fastText for language classification, heuristic rules from MassiveText and C4 for quality filtering, rules- and classifier-based toxicity filtering, and URL, document and paragraph-level deduplication using a Bloom filter.
@@ 3 Building FineWeb
Our design of FineWeb is primarily empirical: we performed a series of “data ablation” experiments to test different methods at each stage of the pipeline. In this section, we chronicle our experimental results and design choices. All ablations follow our iterative dataset building process, i.e., the baseline model for each subsection includes only the processing steps from the previous subsections, unless explicitly mentioned. Our results are fully reproducible with code in our datatrove repository.
@@ 3.1 Experimental setup
All training was performed using the nanotron library. Data ablation models all had 1.71B parameters (including embeddings), used the Llama architecture [2] with a sequence length of 2048, a global batch size of ~2 million tokens, and the GPT2 tokenizer [20]. Within a given experiment, all models were trained on the same amount of data for the same number of steps. Filtering ablations were trained on ~28 billion tokens (roughly the Chinchilla-optimal training size for this model size [1]), while some deduplication ablations and runs to confirm cumulative relative performance improvements after each step of filtering were conducted on 350 billion tokens. The full training hyperparameter are available in Appendix D. We make all models trained for our ablations publicly available on our dataset repository. In total, we trained over 70 models on our internal cluster, for an estimated total of 80,000 H100 GPU hours.
@@ 3.2 Text extraction
Common Crawl data is available in two different formats: WARC and WET. WARC (Web ARChive format) files contain the raw data from the crawl, including the full page HTML and request metadata. WET (WARC Encapsulated Text) files provide a text-only version of crawled websites by using htmlparser [48]. While WET files are commonly used as a starting point for dataset creation, similarly to Gao et al. [29], we found that WET files retained too much boilerplate and menu text. We therefore experimented with extracting the text content from the WARC files using the open source trafilatura library [49], which from visual inspection of the results provided good quality extraction when compared to other available libraries (less boilerplate and menu text). Custom text extraction is relatively costly, but its effects are felt on model performance: Fig. 2 shows the performance of ablation models trained on either trafilatura applied to WARC data or WET data, with minimal additional filtering (fastText language identification to filter samples with English as the highest probability label) and no deduplication. Using trafilatura-extracted text from WARC files clearly results in a more performant model and we therefore use WARC-based data in all of our following experiments.
Figure 1: Trafilatura-extracted WARC vs WET 28B tokens ablation study. Custom text extraction outperforms the default WET data. No filtering or deduplication was applied except fastText English language filtering.
@@ 3.4 Deduplication
Following RefinedWeb [50], we experimented with MinHash, a fuzzy hash-based deduplication technique that scales efficiently to many CPU nodes and allows tuning of similarity thresholds (by controlling the number and the number of hashes per bucket) as well as the length of the subsequences considered (by controlling the n-gram size). We chose to collect each document’s 5-grams, obtained using an English word tokenizer [57], and computed MinHashes using 112 hash functions in total, split into 14 buckets of 8 hashes each — targeting documents that are at least 75% similar. Documents with the same 8 MinHashes in any bucket are considered duplicates of each other. We then perform a transitive clustering step where documents A, B and C will be in the same duplicate cluster if A and C are duplicates and B and C are duplicates, even if A and B do not have 8 matching MinHashes in any bucket with each other. One (randomly chosen) document is kept per duplicate cluster while the remaining duplicates are removed. We further discuss deduplication parameters in Section E.1.
@@
To attempt to improve upon individual deduplication of each snapshot, we additionally experimented with “lighter” global deduplication techniques. Ultimately, none of these techniques improved performance over independent per-snapshot deduplication. A full description of these methods and results are in Section E.3.
@@ 3.6 Developing additional heuristic filters
Past work has mainly developed heuristic filters through data inspection [15]. In this work we devised a more systematic process for designing heuristic filters and tuning their thresholds. We started by collecting over 50 high-level statistics ranging from document-level metrics (e.g. number of lines, avg. line/word length, etc) to inter-document repetition metrics (inspired by MassiveText [1]) on both a high- and low-quality web dataset. Specifically, we used the individually and globally deduplicated versions of the 2013-48 snapshot (previously mentioned in Section 3.4) as our “high-quality” and “low-quality” datasets respectively. We then identified metrics for which the distribution of values differed significantly across the two datasets, inspected the histograms of the two distributions and empirically chose thresholds that would target sections of the histogram where the lower quality dataset frequency was higher than on the corresponding higher quality dataset section. As an example, we plot the distribution of the fraction of lines ending with punctuation metric in  Fig. 8. We can see that the higher quality dataset has in general higher document density for larger values of our metric, and, in particular, the lower quality dataset has a much higher density of documents for values < 0.12. We thus conclude that documents with a fraction of lines ending with punctuation < 0.12 are generally lower quality and use this value as a tentative threshold to filter documents. Following this process for all metrics yielded 16 candidate metric-threshold pairs.
@@
We then assessed the effectiveness of these 16 newly created filters by conducting several 28B token ablation runs on the 2019-18 crawl. Full details for these runs are in  Section E.4. Out of all those runs, we identified three filters (see their ablations runs in  Fig. 7) that demonstrated the most significant improvements on the aggregate benchmark score. Specifically, the chosen filters remove documents where the fraction of lines ending with punctuation is $<=0.12$ (10.14% of tokens removed vs. 30% from the original C4 terminal punctuation filter), where the fraction of characters in duplicated lines is $>=0.1$ (12.47% of tokens removed; the original MassiveText threshold for this ratio is $>=0.2$), and/or where the fraction of lines shorter than 30 characters is $>=0.67$ (3.73% of tokens removed). When applying the three together, ~22% of tokens were removed and the aggregate score increased by about 1% in the 28B token ablations. These filters allowed us to further improve performance and, notably, surpass the C4 dataset performance while filtering out a smaller proportion of data.
@@ 3.7 The final FineWeb dataset
In Fig. 10 we compare FineWeb with the following commonly used openly accessible web-scale datasets: RefinedWeb (500B tokens) [50], C4 (172B tokens) [15]; the Common Crawl-based part of Dolma 1.6 (3T tokens) and 1.7 (1.2T tokens) [58], The Pile (340B tokens) [29], SlimPajama (627B tokens) [35], the deduplicated variant of RedPajama2111RedPajama2 includes 40+ quality annotations but is only actually filtered with the CCNet pipeline[28] (20T tokens) [34], English CommonCrawl section of Matrix (1.3T tokens) [59], English CC-100 (70B tokens) [60], and Colossal-OSCAR (850B tokens) [61]. Notably, FineWeb shows strong performance and FineWeb-Edu (detailed below) outperforms all other open datasets we compared on our aggregate group of tasks, further validating the design choices we made. Note that to train these models we randomly sampled 350 billion tokens from each dataset, without upsampling any individual Common Crawl snapshot.
@@ 4 FineWeb-Edu
To build the synthetic annotations, we use Llama-3-70B-Instruct to score 460,000 randomly sampled webpages from the FineWeb CC-MAIN-2024-10 snapshot for their educational quality on a scale from 0 to 5. We explored several prompt formats to automatically extract an educational score using an LLM and found that the additive scale used in previous work Yuan et al. [63] worked best. It allows the LLM to evaluate each criterion and build the score step-by-step, unlike the single-rating scale [64] which assigns a fixed score based on predefined categories. To avoid having the LLM favor highly technical pages like arXiv abstracts and submissions, we prompted it to focus on grade-school and middle-school level knowledge. The prompt used for synthetic annotations is in Section F.1.
@@
To confirm the effectiveness of education filtering at a larger scale, we conducted a larger ablation training a 1.71B model on 350 billion tokens, similar to the FineWeb filtering ablations mentioned above. As shown in Fig. 10 and Fig. 11, we observed that FineWeb-Edu surpasses FineWeb and all other open web datasets, with quite remarkable improvements on educational benchmarks such as MMLU, ARC and OpenBookQA. Specifically, MMLU score increases from 33% to 37%, a relative improvement of approximately 12%, and ARC score goes from 46% to 57%, an improvement of about 24%.
On MMLU, FineWeb-Edu can match the final performance of Matrix with almost 10x fewer tokens, demonstrating the effectiveness of classifiers trained on LLM annotations for large-scale data filtering. Additional evaluation plots can be found in Section F.2.
4.1 Topic distribution
To examine how the educational classifier may skew the dataset towards certain topics, we embed [67] 50k samples from FineWeb and 50k samples from FineWeb-Edu using a sentence-transformers [68] model (all-MiniLM-L6-v2) which we then project to 2D space using UMAP [69]. Finally, we use DBSCAN [70] clustering to find the 100 densest topic clusters in the union of the two datasets, which we label using Llama 3.1 70B [6]. To compare the two datasets, we plot the difference of the size (as a percentage of the entire dataset) of each cluster in FineWeb-Edu and FineWeb in Fig. 18. The educational classifier heavily favors topics such as ’Education, Learning, Teaching’ or ’History, Culture, Politics’, while down-sampling ’Business, Finance, Law’, ’Entertainment, Film, Theater’ and ’Places, Travel, Real Estate’, among others.
4.2 Domain fit
We evaluate the macro average perplexity of six checkpoints from our FineWeb and FineWeb-Edu ablation models on the domains from Paloma [71]. We use the codebase provided in [71] but intentionally do not perform decontamination, to compare how well each dataset covers different domains. The results are in Fig. 12. The FineWeb model generally shows lower perplexity in broad web sources such as C4, mC4, Falcon, Dolma V1.5 or the CommonCrawl subset of RedPajama, as well as on Twitter AAE, Manosphere, Gab, reddit (100 Subreddits) or 4chan. FineWeb-Edu tends to favour sources containing Wikipedia (WikiText-103 and M2D2 Wikipedia) or that are heavy in academic content (M2D2 S2ORC, which has semantic scholar papers, and the Arxiv subset of RedPajama). FineWeb-Edu also seems to have better coverage of programming content (100 PLs) than FineWeb. We have included results per subset for some of these sources in Section F.4.
Figure 12: FineWeb and FineWeb-Edu fit to Paloma domains. FineWeb has lower perplexity on broad web sources while FineWeb-Edu has better coverage of Wikipedia and programming content.
@@ 5 Bias analyses
Language models are known to reflect the biases present in their pretraining datasts [72, 73, 74, 75, 76, 77]. To provide a brief picture of dataset bias in FineWeb and FineWeb-Edu, we focus on subgroups recognised as “sensitive” or “protected” in English-speaking countries. These are a subset of subgroups that are historically subject to discrimination and are disproportionately the target of negative societal norms such as stereotyping, which is reflected in text-based language consumed for a dataset. We find that the FineWeb dataset has a relative overrepresentation of words that reflect hegemonic norms, known to be overrepresented in online text [72], such as ‘man’ and ‘christian’. Although biases across the gender, religion, and age subgroups examined are not strong, we see the most skewed association between religion words and intimacy, such as ‘christian dating’ and ‘jewish singles’. Fittingly, the FineWeb-Edu dataset captures associations that are less tied to intimacy compared to FineWeb and more expected from educational content of history and health, such as ‘man’ being associated to ‘king’, and ‘woman’ associated to ‘pregnancy’. Further details are provided in Appendix G.
@@ Appendix C Linked resources
FineWeb Llama3 annotations https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu-llama3-annotations
Educational classifier https://huggingface.co/HuggingFaceFW/fineweb-edu-classifier
@@
Appendix D Data ablation setup
D.1 Model architecture
Parameter Value
Architecture Llama
Number of attention heads 32
Number of hidden layers 24
Number of key-value heads 32
RMS Norm epsilon 1e-05
Tied word embeddings True
Embedding size 50257
Total number of parameters 1.71B
Random initialization std 0.02
Tokenizer GPT2
D.2 Distributed training setup
Parameter Value
Data parallelism (dp) 64
Tensor parallelism (tp) 1
Pipeline parallelism (pp) 1
Micro-batch size 4
Sequence length 2048
Batch accumulation per replica 4
D.3 Optimizer Configuration
Parameter Value
Adam beta1 0.9
Adam beta2 0.95
Adam epsilon 1.0e-8
Gradient clipping 1.0
Weight decay 0.1
Learning rate 3e-4
Warmup steps 500
Warmup style linear
Decay style cosine
Minimum decay LR 3.0e-5
Appendix E Deduplication
E.1 Deduplication parameters
@@
E.2 Measuring the effect of deduplication
@@
E.3 Alternative global deduplication